OpenStudy (anonymous):

Approximately 4.5 liters of blood circulates in the body of the average human adult. How many quarts of blood does the average person have? Use the conversion equivalency 1 gallon = 3.785 liters and 4 quarts = 1 gallon.

OpenStudy (perl):

$$ \Large 4.5~ liters \cdot \frac{1~ gallon}{3.785~ liters} \cdot \frac{4 ~ quarts}{1~ gallon} $$

OpenStudy (perl):

do you see how the units cancel